John accepted a reduction of 15% in his salary when his company was not doing well. Now his company's financial position has improved and his boss wants to restore his original salary By what percentage must his reduced salary be increased?
Answer by jim_thompson5910(35256)

x - 15% of x = x - 0.15x = 0.85x

-------------------------------------------------------

Old Salary: x dollars

New Salary: 0.85x dollars

-------------------------------------------------------

Now we want to go from 0.85x to x dollars again.

Difference: x - 0.85x = 0.15x

Divide this difference by the amount you are increasing from (which is 0.85x)

0.15x/0.85x = 0.17647

0.17647 ----> 0.17647*100 = 17.647%

So you need to increase his salary by roughly 17.647% to get back to the original salary.

Side Note: a common mistake is to think that you increase it by 15% to get back to the original, but taking 15% of a smaller salary leads to a smaller increase. So that's why the percentage needs to be bumped up to get it back to the original.